OpenStudy (anonymous):

Composition of functions help! f(x)=-3x+2 g(x)=3x^2-x+2 Find (f+g)(x)

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

all I know so far is to start off with g(f(x)) I think

OpenStudy (anonymous):

.............

OpenStudy (jhannybean):

Add the two functions together

OpenStudy (anonymous):

so (-3x+2)+(3x^2-x+2)

OpenStudy (jhannybean):

Yes.

OpenStudy (anonymous):

x-x+4?

OpenStudy (jhannybean):

\[(f+g)(x) = (-3x+2)+(3x^2-x-2)\]\[3x^2 -x-3x+2-2 = ?\]

OpenStudy (anonymous):

3x^2-4x

OpenStudy (jhannybean):

Good job

OpenStudy (anonymous):

thats it?

OpenStudy (jhannybean):

You can also factor our an x, since it's common between both terms.\[\large 3x^2-4x = x(3x-4)\]

OpenStudy (jhannybean):

But that's if your teacher wants it in fully factored form,otherwise you're good to go :)

OpenStudy (anonymous):

really? that was much simpler than I thought it's be

OpenStudy (jhannybean):

Yeah, haha.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (jhannybean):

np

OpenStudy (anonymous):

I need your help with another problem like this